OpenStudy (anonymous):

How do I divide polynomials? I am trying to take a test, but the online textbook never went over dividing polynomials, and I need help. How do I find the degree of the polynomial that results from division?

OpenStudy (anonymous):

And this is not a direct question. This is just to assist me.

OpenStudy (anonymous):

\[\frac{ x ^{2} }{ x ^{3} } = x ^{2-3}\]

OpenStudy (anonymous):

How do I divide a binomial by a trinomial?

OpenStudy (dtan5457):

You can factor the trinomial and cross out like terms

OpenStudy (anonymous):

@PretzelEater24 Reported

OpenStudy (anonymous):

Pretzel, you should realize the circumstances I'm in. I actually have been looking at this for over an hour and cannot seem to find the way to do so.

OpenStudy (anonymous):

Not everyone here just suddenly asks questions for direct answers.

OpenStudy (anonymous):

Sorry.

OpenStudy (anonymous):

@MathematicsNerd don't mind him, just feeding is a-hole behavior.

OpenStudy (dtan5457):

Agreed, I'm sure he asks plenty of questions himself. He's not a genius, and with that hypocritical behavior, he never will be

OpenStudy (anonymous):

@dtan5457 has it right btw, factor the trinomial then cross out like terms.

OpenStudy (anonymous):

@Sadworld I'm sorry.

OpenStudy (anonymous):

(x^2+2xy+y^2)/(x+y) = (x+y)(x+y)/(x+y) = (x+y)

OpenStudy (anonymous):